OpenStudy (anonymous):

what is the midpoint of (-19,0) and (-4,4)

OpenStudy (kobeni-chan):

(-11,5, 2)

OpenStudy (kobeni-chan):

*-11.5

OpenStudy (anonymous):

How do u find that?

OpenStudy (kobeni-chan):

You find the average of the x and y variables. -19 + -4 = -23 -23/2 = -11.5 0 + 4 = 4 4/2 = 2

OpenStudy (anonymous):

Oh! Thanks i have always wondered this :P